### v1

# Large Language Models as Attribution Regularizers for Efficient Model Training

Abstract
Large Language Models (LLMs) have demonstrated remarkable performance across diverse domains. However, effectively leveraging their vast knowledge for training smaller downstream models remains an open challenge, especially in domains like tabular data learning, where simpler models are often preferred due to interpretability and efficiency.
In this paper, we introduce a novel yet straightforward method for incorporating LLM-generated global task feature attributions into the training process of smaller networks. Specifically, we propose an attribution-matching regularization term that aligns the training dynamics of the smaller model with the insights provided by the LLM. By doing so, our approach yields superior performance in few-shot learning scenarios. Notably, our method requires only black-box API access to the LLM, making it easy to integrate into existing training pipelines with minimal computational overhead.
Furthermore, we demonstrate how this method can be used to address common issues in real-world datasets, such as skewness and bias. By integrating high-level knowledge from LLMs, our approach improves generalization, even when training data is limited or imbalanced. We validate its effectiveness through extensive experiments across multiple tasks, demonstrating improved learning efficiency and model robustness.
Keywords:
Large Language Models Attribution Regularization Data-Efficient Learning1 Introduction
The recent expansion in model parameters and training data for large language models (LLMs) has driven a significant breakthrough in natural language processing (NLP) [8, 12, 24, 49]. These models exhibit remarkable performance across various evaluation paradigms, such as zero-shot [32] and few-shot inference, leveraging in-context learning [36, 51]. This capability stems from the extensive text corpora used for training, which embed rich prior knowledge into LLMs, allowing them to approximate expert knowledge across diverse domains. Although their strong performance and generalization capabilities have been successfully extended to other modalities, such as images [34] and speech [54], their application in tabular learning settings remains limited.
Several challenges hinder the adoption of LLMs for tabular data tasks. Firstly, their large parameter counts demand substantial computational resources, typically reliant on GPUs, which significantly increases operational costs. Secondly, tabular learning is often employed in domains where transparency and interpretability are critical, such as healthcare and finance. In these fields, simpler and more interpretable models, such as logistic regression or decision trees, are often preferred. Although techniques exist to enhance LLM interpretability [3], they currently fall short compared to the inherent explainability of simpler models.
In this paper, we investigate the potential of utilizing LLMs as training regularizers to enhance few-shot learning performance and improve generalization, especially in scenarios with skewed or biased training data. Specifically, we propose a method called Large Language Model Attribution Aligned Training (LAAT), which introduces an attribution-matching regularization term that aligns the local, feature-wise explanations of smaller models with the global, task-specific explanations generated by LLMs. This approach harnesses the strong generalization capabilities of LLMs while preserving the efficiency and transparency of smaller models. We share our code at: https://github.com/davor10105/laat
2 Related Work
2.1 Standard Machine Learning Approaches
Inspired by the success of deep learning in other domains, numerous efforts have sought to apply self-supervised learning to tabular data to develop transfer learning-ready models. These approaches include masked feature prediction [4, 35], feature corruption correction [6, 53], and contrastive pre-training [48]. However, comparative studies indicate that gradient-boosted tree ensembles still outperform these methods [20, 45]. More recently, Nam et al. [38] introduced Self-generated Tasks from UNlabeled Tables (STUNT), leveraging self-generated few-shot tasks for tabular learning, though its reliance on large unlabeled datasets may limit practical applicability. Additionally, Hollmann et al. [25] proposed the Tabular Prior-data Fitted Network (TabPFN), a tabular foundation model pre-trained on millions of synthetic datasets.
2.2 Large Language Models in Tabular Learning
Most approaches integrating large language models (LLMs) into tabular learning rely on encoding task and feature descriptions in natural language, serializing the data, and leveraging LLMs for inference—either through in-context learning [46] or additional fine-tuning [13, 23, 50]. However, these methods face significant drawbacks, including the high cost of LLM inference for individual samples and the computational demands of fine-tuning.
In sensitive domains such as medicine or finance [44], where transparency is critical, the opaque decision-making of LLMs is less desirable than traditional, smaller models. Alternative approaches involve using LLMs to generate synthetic examples to augment existing datasets, employing both in-context learning and fine-tuning [42, 47, 55]. However, these methods inherit the same scalability issues, particularly when dealing with high-dimensional datasets, where generating sufficiently large datasets becomes computationally expensive.
Recently, Han et al. [21] introduced FeatLLM, a novel approach that utilizes LLMs as feature engineers. Instead of directly performing inference, FeatLLM employs code-generating LLMs to create preprocessing functions that transform the original dataset into a more suitable representation for few-shot classification. This method implements an ensemble classifier to combine insights from multiple feature transformations, improving robustness and classification accuracy. FeatLLM significantly reduces resource requirements by relying solely on pre-trained LLMs with API-level access. Moreover, FeatLLM outperforms existing fine-tuned and in-context learning approaches while maintaining lower computational costs.
2.3 Explanation Guided Learning
A growing line of research explores enhancing model behavior through additional supervision derived from explainable artificial intelligence (XAI) techniques. This field can be broadly categorized into local explanation-guided learning and global explanation-guided learning [17].
Local explanation guidance applies supervision signals or regularization terms to individual model explanations, steering learning at the sample level. This approach is more prevalent due to the extensive development of local explanation techniques, particularly in the image domain, such as Grad-CAM [43], Layer-wise Relevance Propagation (LRP) [5], and attention-based attributions [2]. Ross et al. [41] propose regularizing differentiable models by penalizing input gradients, aligning them with expert-defined attribution maps. Dharma et al. [29] use object bounding boxes as explanation supervision signals. In text classification, several studies leverage per-sample human-annotated rationales [11, 28, 56]. Gao et al. [18] demonstrate the effectiveness of local explanation supervision under limited training data. However, a key limitation of this approach is its reliance on per-sample attribution annotations, which are often difficult and costly to obtain, particularly in expert-driven fields like medicine.
Global explanation guidance, in contrast, does not require instance-level attributions, instead offering a broader, more scalable approach to shaping model behavior. Liu et al. [33] reduce undesired biases by penalizing nonzero attributions on sensitive tokens. Erion et al. [16] aggregate local feature attributions via expected gradients to improve interpretability. Weinberger et al. [52] extract prior knowledge from multiple gene expression datasets to construct meta-features, training a deep global attribution model alongside a predictive model with a regularization loss. However, this method assumes the availability of additional datasets related to the problem, which may not always be feasible.
3 Method
In contrast to other methods that utilize LLMs for tabular data prediction, we seek to minimize both the computational and price overhead of their use, while simultaneously still effectively using their generalization abilities and providing small, interpretable models that can be readily used in existing pipelines.
3.1 Formulation
Given a trained binary classification model parametrized by , an attribution produced by an attribution method for an input is a vector , where is the attribution score of the input feature . We are interested in the expected value of the attribution scores over the entire dataset, given by , where represents the data distribution.
For certain datasets where the expected attribution follows an intuitive pattern that humans can interpret, we hypothesize that this expected value can be approximated using a large language model and thus serve as a valuable local attribution guide during training.
Taking a step back, given an untrained model along with a task description and feature descriptions, we query an LLM to generate importance scores for each feature, producing a vector . During model training, we then regularize the local attribution scores of the model to align with these LLM-derived scores. This regularization acts as a guiding signal, helping the model maintain behavior that aligns with intuitive, human-understandable reasoning.
The final model’s loss function consists of two components: the standard binary cross-entropy loss and an attribution regularization term. The regularization term is the mean squared error between the normalized attribution scores and the normalized LLM-derived scores, weighted by . The overall loss is given as a weighted sum of these terms:
Following Ross et al. [41], we employ the input gradient as our chosen attribution method.
3.2 LLM Prompting and Scores Parsing
To enable Large Language Models (LLMs) to generate meaningful feature attribution scores that effectively guide downstream models during training, we developed a comprehensive prompting methodology. This approach incorporates several key components to maximize score accuracy and relevance.
3.2.1 Task and Dataset Contextualization
We first embed relevant task and dataset information within the prompt structure by incorporating key details. Task descriptions provide a concise overview of the classification objective and possible outcomes. For example: "Predict whether this patient’s breast cancer will reoccur. Yes or no?" — an approach aligned with established methodologies in recent literature [21, 23]. Feature descriptions ensure clarity by defining each dataset attribute; for instance, "Age: The age of the patient at the time of diagnosis." To handle categorical features, we apply one-hot encoding and include specific descriptions for each category. This allows the LLM to generate distinct attribution scores for individual categorical levels rather than the overall feature.
3.2.2 Score Generation Protocol
We implement a structured scoring protocol by instructing The LLM to assign an integer score between and for each feature, creating a standardized scale of feature importance. Following chain-of-thought methodologies [51], we incorporate prompts that require explicit reasoning before score assignment. This approach has demonstrated superior performance in complex reasoning tasks and improves the interpretability of the scoring process. The complete prompt template used for score generation is provided in Appendix 1
3.2.3 Score Extraction and Aggregation
To transform the LLM’s textual responses into structured numerical data we employ a secondary LLM instance configured for function calling to extract numerical scores from the primary LLM’s reasoning and output. This parser converts the semi-structured scores into a standardized list format suitable for downstream processing.
To mitigate potential variability in LLM outputs, we generate scores multiple times () and compute their mean. This ensemble approach produces the final LLM-provided feature-wise attribution vector (), enhancing the stability and reliability of the attribution signal.
This methodology creates a robust framework for leveraging LLM capabilities to generate informative feature attribution scores that can effectively guide downstream predictive models during their training process.
4 Experiments
We conducted a comprehensive evaluation of LAAT across diverse tabular datasets, examining its performance in few-shot learning contexts and scenarios where significant bias was present in the training data. Furthermore, we performed supplementary experiments to investigate the impact of various hyperparameter configurations on the proposed methodology.
4.1 Few-shot Learning
While large volumes of data are readily available in many domains, expert-labeled data remains scarce in fields requiring specialized knowledge, such as medicine. To address this challenge, recent research has focused on enhancing the generalization capabilities of tabular classification models under minimal labeled data constraints [21, 23, 38]. In this experiment, we assess the effectiveness of our proposed approach in leveraging LLM-derived knowledge as a guiding signal during extremely low-shot training scenarios.
We evaluate our approach on seven publicly available binary classification datasets:
adult [7] - predicting whether an individual earns over $50,000 annually
bank [37] - predicting whether a client will subscribe to a term deposit
breast-ljub [57] - predicting whether a patient’s breast cancer will reoccur
cdc-diabetes [1] - predicting whether an individual has diabetes
diabetes [27] - predicting whether an individual has diabetes
electricity [22] - predicting the price change of electricity (up or down) in New South Wales
myocardial [19] - predicting whether the myocardial infarction complications data for an individual shows chronic heart failure
The datasets vary in size and complexity, and their additional basic information is outlined in Appendix 5.
4.1.1 Baselines
4.1.2 Implementation Details
We employ three distinct foundation models as our importance score estimators: Llama 3.3 70B () [15], Gemini 2.0 Flash () [40] and GPT-4o-mini () [26]. As simple downstream models, we utilized logistic regression and a two-layer multilayer perceptron (MLP) with ReLU activation and 100 hidden units. Optimization was performed using the Adam optimizer [30] with a learning rate of and no weight decay, leveraging LAAT’s internal regularization via the factor, which was set to 100. Additionally, the number of importance score estimations for LAAT was set to 5. LAAT was executed without early stopping. Data preprocessing involved one-hot encoding categorical variables and standardizing numerical features using z-score normalization. Exact task descriptions used can be found in Appendix 6.
For traditional machine learning models (logistic regression, MLP, random forest, XGBoost, and CatBoost), we performed hyperparameter optimization through grid search with 5-fold cross-validation, with the exception of 1-shot learning scenarios where default scikit-learn [39] hyperparameters were applied. The complete baseline hyperparameter search spaces are detailed in Appendix 7. TabPFN and FeatLLM were implemented with their respective default parameter configurations as described in the original publications. Additionally, FeatLLM incorporated early stopping mechanisms in the 5-shot and 10-shot experimental settings. Due to the inherent complexity of generating multiple conditions and preprocessing functions in the FeatLLM method, Llama 3.3 exhibited inconsistencies in producing valid outputs. As a result, we excluded it from the FeatLLM experiments.
To ensure statistical robustness, we conducted ten independent experimental runs for each model.
4.1.3 Results
The results of our experiments, averaged over ten runs, are presented in Tables 1 (for non-LLM methods) and 2 (for LLM-based methods). As shown, LAAT models consistently rank among the top-performing methods, achieving 16 top scores across 21 experiments. Notably, the highest-performing variants of LAAT were and , each securing 6 top scores. Furthermore, the LAAT-trained models outperformed their non-LAAT counterparts in 15, 19, and 20 out of 21 experiments using logistic regression with Llama 3.3 70B, Gemini 2.0 Flash, and GPT-4o-mini as scoring models, respectively. For MLP, LAAT models exceeded the performance of vanilla MLP in 17, 20, and 20 experiments, respectively. These results underscore the positive impact of LLM attribution alignment on the generalization capabilities of even simple models, achieving significant improvements without the need for extensive hyperparameter tuning, which is often required for other approaches.
Dataset | Shot | LR | MLP | RF | XGB | CatBoost | TabPFN |
adult | 1 | ||||||
5 | |||||||
10 | |||||||
bank | 1 | ||||||
5 | |||||||
10 | |||||||
breast-ljub | 1 | ||||||
5 | |||||||
10 | |||||||
cdc-diabetes | 1 | ||||||
5 | |||||||
10 | |||||||
diabetes | 1 | ||||||
5 | |||||||
10 | |||||||
electricity | 1 | ||||||
5 | |||||||
10 | |||||||
myocardial | 1 | ||||||
5 | |||||||
10 |
Method | FeatLLM | LAAT | |||||||
Model | Ensemble | LR | MLP | ||||||
Dataset | Shot | ||||||||
adult | 1 | ||||||||
5 | |||||||||
10 | |||||||||
bank | 1 | ||||||||
5 | |||||||||
10 | |||||||||
breast-ljub | 1 | ||||||||
5 | |||||||||
10 | |||||||||
cdc-diabetes | 1 | ||||||||
5 | |||||||||
10 | |||||||||
diabetes | 1 | ||||||||
5 | |||||||||
10 | |||||||||
electricity | 1 | ||||||||
5 | |||||||||
10 | |||||||||
myocardial | 1 | ||||||||
5 | |||||||||
10 |
Figure 1 presents the average ROC AUC performance across all datasets. While FeatLLM achieves the highest score in the 1-shot scenario, LAAT models follow closely, surpassing all non-LLM baselines. However, as FeatLLM’s performance plateaus beyond the initial 1-shot setting, LAAT models demonstrate a substantial advantage in both the 5-shot and 10-shot scenarios, outperforming all other methods by a significant margin. Among the scoring models evaluated, GPT-4o-mini emerged as the top performer, with Gemini 2.0 Flash-Lite following closely behind. Although Llama 3.3 70B exhibited lower performance compared to these two LLMs, it still outperformed all non-LLM baselines in the 1-shot setting, surpassed all baselines in the 5-shot setting, and maintained competitive performance in the 10-shot experiments.
4.2 Learning on Biased Data
Many real-world datasets exhibit varying degrees of bias, which can negatively impact machine learning models by introducing misleading correlations. For instance, a model trained on historical recruitment data may inherit biases related to gender and job roles [10]. Similarly, systemic biases in criminal justice datasets often lead to strong correlations between demographic attributes and outcomes, potentially reflecting enforcement patterns rather than actual crime rates. In some cases, these confounding factors can be explicitly removed when they are demonstrably unrelated to the target classification. However, the challenge becomes more complex when the relationship is "soft"—i.e., a correlation exists, but its exact influence is difficult to quantify.
To evaluate the performance of baseline methods alongside our proposed LAAT method, we introduce artificial biases into the training portion of the medicine-related datasets used in the previous experiment while keeping the evaluation data distribution unchanged. This approach mimics real-world biases that may arise due to the way data is collected—for example, demographic skews in clinical trials or socioeconomic biases in electronic health records. By training models on biased distributions and assessing them on the original, unbiased test data, we simulate practical scenarios where spurious correlations may mislead models, allowing us to analyze their robustness to such biases. Specifically, we apply the following modifications:
breast-ljub - Exclude all patients under the age of 50 who experienced breast cancer recurrence, introducing a bias that overestimates recurrence in elderly patients.
cdc-diabetes - Exclude all women with diabetes and all men without diabetes, artificially inflating the apparent prevalence of diabetes among men.
diabetes - Remove all individuals under 50 years old who have diabetes, encouraging models to over-rely on age for diabetes prediction.
myocardial - Remove all female patients with a history of myocardial infarction, biasing models toward overestimating the risk of myocardial infarction in men.
4.2.1 Results
As in the previous experiment, results are presented in Tables 3 for non-LLM methods and 4 for LLM-based methods. LAAT effectively mitigates dataset biases and achieves the highest performance across all datasets, surpassing non-LLM baselines by more than 20 ROC AUC points in some cases. While FeatLLM outperforms traditional baselines, it remains significantly behind LAAT models. The superior performance of LAAT on this challenging benchmark further highlights the benefits of attribution guidance — not only in enhancing performance in data-scarce scenarios but also in addressing subtle, hard-to-detect dataset biases that could otherwise compromise the integrity of the machine learning pipeline.
Model | LR | MLP | RF | XGB | CatBoost | TabPFN |
Dataset | ||||||
breast-ljub | ||||||
cdc-diabetes | ||||||
diabetes | ||||||
myocardial |
Method | FeatLLM | LAAT | ||||||
Model | Ensemble | LR | MLP | |||||
Dataset | ||||||||
breast-ljub | ||||||||
cdc-diabetes | ||||||||
diabetes | ||||||||
myocardial |
4.3 Hyperparameter Analysis
We investigate the impact of varying the regularization factor and the number of importance score estimations on the final performance of LAAT models. Specifically, we assess the performance of a logistic regression model trained using LAAT alignment with importance scores provided by GPT-4o mini. The experimental results are presented in Figures 3 and 3, corresponding to the variation and the number of importance score estimations, respectively.
Our findings indicate that increasing initially enhances model performance, reaching an optimal value around . However, further increases beyond lead to a substantial decline in performance. In contrast, increasing the number of importance score estimations results in a steady performance improvement, which plateaus at approximately four to five estimations, depending on the k-shot setting. These results suggest that while additional score generation and ensembling can enhance performance, their benefits diminish beyond a certain threshold.
5 Conclusion
We introduce Large Language Model Attribution Aligned Training (LAAT), a novel approach that leverages importance scores inferred by large language models (LLMs) as local attribution guides. This method effectively harnesses the generalization capabilities of LLMs while employing simple, traditional machine learning models. Although gradient boosting trees have traditionally dominated tabular data learning, LAAT significantly outperforms existing methods in few-shot learning and biased dataset scenarios. Notably, it surpasses FeatLLM, the current state-of-the-art LLM-based few-shot learning approach.
Beyond its standalone effectiveness, LAAT is highly versatile and can be used in conjunction with more complex models, provided they are differentiable — for example, TabPFN. Additionally, it can be combined with feature preprocessing techniques such as FeatLLM. Given the interpretability of LLM-derived importance scores, LAAT could be incorporated into interactive, chat-based interfaces, enabling human experts to refine these scores before finalizing them. This human-in-the-loop refinement could further enhance model performance by leveraging domain expertise. Furthermore, LAAT’s framework could be extended beyond tabular data to other modalities, such as images, by integrating it into concept bottleneck models [31]. This extension would allow LAAT to guide feature selection and attribution in vision tasks, improving interpretability and generalization in image-based applications. We leave the exploration of these directions to future work.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix
We provide additional dataset information, task descriptions, score extraction prompt and hyperparameter search spaces utilized in our experiments.
Basic Dataset Information
Dataset | Samples | Num. features | Cat. features | Pos. label ratio |
adult | 48842 | 14 | 8 | 23.93% |
bank | 45211 | 16 | 9 | 11.70% |
breast-ljub | 277 | 16 | 0 | 29.24% |
cdc-diabetes | 253680 | 21 | 0 | 13.93% |
diabetes | 768 | 8 | 0 | 34.90% |
electricity | 45312 | 8 | 0 | 42.45% |
myocardial | 686 | 91 | 13 | 22.16% |
Task Descriptions
Dataset | Task Description |
adult | Predict whether this person earns more than 50000 dollars per year. Yes or no? |
bank | Predict whether this client will subscribe to a term deposit. Yes or no? |
breast-ljub | Predict whether this patient’s breast cancer will reoccur. Yes or no? |
cdc-diabetes | Predict whether the patient has diabetes. Yes or no? |
diabetes | Predict whether the patient has diabetes. Yes or no? |
electricity | In the electricity market, prices are not fixed and are affected by demand and supply of the market. They are set every five minutes. Electricity transfers in the state A to/from the neighboring state B are done to alleviate fluctuations. Based on the current measurement, predict whether the price of electricity in state A will go up. Yes or no? |
myocardial | Predict whether the myocardial infarction complications data of this patient show chronic heart failure. Yes or no? |
Prompt Template for Importance Score Generation
Hyperparameter Search Spaces
Model | Hyperparameter search space |
Logistic Regression | C: [100, 10, 1, 1e-1, 1e-2, 1e-3, 1e-4, 1e-5] |
MLP |
alpha: [0.001, 0.01, 0.1, 1, 10]
learning_rate_init: [0.1, 0.01, 0.001, 0.0001] |
XGBoost |
max_depth: [2, 4, 6, 8, 10]
alpha: [1e-4, 1e-3, 1e-2, 1e-1, 1, 10] lambda: [1e-4, 1e-3, 1e-2, 1e-1, 1, 10] eta: [0.01, 0.03, 0.1, 0.3] |
Random Forest |
bootstrap: [True, False]
max_depth: [2, 4, 6, 8, 10] n_estimators: [2, 4, 8, 16, 32, 64] |
CatBoost |
colsample_bylevel: [0.01, 0.03, 0.06, 0.1]
boosting_type: ["Ordered", "Plain"] depth: [2, 4, 6, 8, 10] |